OpenStudy (anonymous):

For the investment with interest compounded annually, find the final balance and the total interest earned. $70,600 for 3 years at 4.2%

OpenStudy (anonymous):

A. final balance: $70,896.52 interest earned: $296.52 B. final balance: $73,565.20 interest earned: $2965.20 C. final balance: $79,495.60 interest earned: $8895.60 D. final balance: $79,874.45 interest earned: $9274.45

OpenStudy (anonymous):

if A = P (1+r)^t, then the final amount is A = (70000)(1.042)^3

OpenStudy (anonymous):

can you simplify that?